OpenStudy (anonymous):

What is the cosine rational form of 60 degrees?

OpenStudy (paxpolaris):

 cos(60deg)= 1/2

OpenStudy (barrycarter):

This would be a 30-60-90 triangle, so the smaller side is half the length of the hypotenuse, so 1/2

OpenStudy (anonymous):

Cos60 = 1/2

OpenStudy (anonymous):

The cosine ratio in rational form for a 60° angle is? How exactly do I do this? I understand the first example,

OpenStudy (anonymous):

Well, 60 degrees in the radian form is pi/3 why? pi=180 and 180/3 gives you 60 So, cos60 =cos pi/3 = 1/2 by rule.

OpenStudy (anonymous):

So would the tan90 also be 1/2?